OpenStudy (anonymous):

What is the equation of the line that passes through the points (-2, 1) and (1, 10)? A 3x - y = -7 B 3x - y = -5 C 3x - y = 5 D x + 3y = -5 When I do the problem, I come up with 9... but I forget how to put it into an equation.

OpenStudy (anonymous):

First find the slope by using the rise/run formula. m=(1-10)/(-2-1)=(-9)/(-3)=3. Then you have y=3x+c. Plug in one of the points to find c. 1=3(-2)+c 7=c y=3x+7 or 3x-y=-7